7. Why do dominant alleles tend to be advantageous and recessive alleles tend to be disadvantageous in most populations?

8. You are studying a population of parrots and you observe that there are two alleles for wing color. Allele 1 codes for green wings and is dominant over allele 2, which codes for red wings. Allele 2 is at a frequency of 70% of the population, what should be the frequency of allele 1 homozygotes?

9. You are studying a population of chupacabras that has two alleles for spine length (L for long and S for short, with L dominant to S). In a population you surveyed, you find 25% of individuals with the short-spine phenotype. If you were to consider this information in the context of the Hardy-Weinberg theorm, the 25% you have been given represents: -q -2pq -q^2 -2qn -q^2*n

10. Name one form of selection that can generate genotypic diversity.
Answer #1

10. Environmental heterogeneity is a major factor in maintaining and structuring genetic diversity in natural populations at all geographical scales - global, regional, and local.Genetic polymorphism and heterozygosity in nature are structured nonrandomly on a massive scale.

7.Dominance does not determine whether an allele is deleterious, neutral or advantageous. However, selection must operate on genes indirectly through phenotypes, and dominance affects the exposure of alleles in phenotypes, and hence the rate of change in allele frequencies under selection. Deleterious recessive alleles may persist in a population at low frequencies, with most copies carried in heterozygotes, at no cost to those individuals. These rare recessives are the basis for many hereditary genetic disorders.

8. The frequency of dominant allele ( Allele 2 ) is 70% of population = 0.7

So, we know = p+q = 1

If p= 0.7 then q = 0.3

This means male contribute 0.3 of allele 1 and female also contribute 0.3 of allele 1 at the time of zygote .

So 0.3 * 0.3 = 0.09 = 9%

Hence homozygotes of allele 1 is 9% of the population .

Ques 9 is unclear for me.. If possible write it on paper properly and then post it.